# Information That Matters: Exploring Information Needs of People Affected by Algorithmic Decisions

### v1

@@ 5. Discussion
In this section, we discuss the findings of our study regarding information needs and understanding (Section 5.1), stakeholder aims and explanatory stances (Section 5.2), and measures to incorporate affected stakeholders into explanation design (Section 5.3), and provide a list summarizing the implications of our findings (Section 5.4).
5.1. Eliciting affected stakeholders’ information needs and resolving understanding issues
@@
One of XAI’s enduring key questions is: What makes a ”good” explanation? Recent work outlines that, to produce understanding, explanations should be contrastive, selected, and social, while being first and foremost contextual (Miller, 2019), that they should interface directly with people’s causal reasoning (Byrne, 2023) and consider the learner’s aim and prior knowledge (Mueller et al., 2019). These demands can be realized practically by identifying relevant information needs and examining how this information is processed. In the following, we present how the findings of our study support this analysis.
5.1.1. Collecting information needs and using them as reference for explanation design.
The findings described in Section 4.1 show that affected stakeholders’ information needs cover all parts of an ADM system’s ”lifecycle” (Dhanorkar et al., 2021), ranging from technical details over deployment consequences to different dimensions of usage. Our proposed ”XAI Novice Question Bank” (Figure 2) aims to provide a summary of these topics which can be used for future explanation approaches. In empirical settings, presenting participants with collections like the XAI Novice Question Bank and the XAI Question Bank (Liao et al., 2020) can support the ”intelligibility” (Lim and Dey, 2009) of a system and reduce the risks of ”blind spots” , i.e., information that stakeholders cannot designate or locate.
@@
To ascertain which topics are actually deemed relevant in a given use case, however, requires speaking to the people involved. This is demonstrated by the differences in information needs between use cases A and B as well as the differences to Liao et al. (2020)’s study. In line with previous work, we observe that various contextual factors mediate which information need participants articulate, including: familiarity with the presented system and comparisons with similar systems (Kramer et al., 2018), tangibility of the system (Long et al., 2021), perceived risks and benefits (Araujo et al., 2020) (e.g., trade-off between safety and privacy in the wristband use case), domain expertise (Wang and Yin, 2021) (e.g., questions about development), and technical expertise (Cheng et al., 2019).
@@
Our findings highlight two key issues of how on-explanations and understanding interact: First, letting stakeholders acquire information independently supports personalized explanations (Conati et al., 2021), but at the same time transfers to stakeholders the burden of deciding when a sufficient understanding has been reached, which can be difficult to assess (Keil, 2006). In consequence, self-reports of understanding can be unreliable and prone to the ”illusion of explanatory depth” (Rozenblit and Keil, 2002; Chromik et al., 2021): an overestimation of one’s understanding. Previous work has outlined various strategies to offset these issues, including deliberate self-explanation (Duckworth, 2001), better initial introductions to the system (Chromik et al., 2021), and actively checking for people’s understanding by letting them predict system outcomes before providing explanations (Buçinca et al., 2020). We argue that further research is needed to determine suitable formats for self-assessments of explanatory understanding in XAI.
@@
Second, several issues can interfere with information acquisition, including blind spots and outsourcing. We use the term blind spots for information that participants missed but could not articulate, which might correspond to a lack of ”intelligibility types” (Lim and Dey, 2009), i.e., a lack of knowledge about available information and ways to inquire it. References like our proposed XAI Novice Question Bank (Figure 2) and the XAI Question Bank (Liao et al., 2020) should help in addressing this issue by outlining the relevant categories of information. Outsourcing, on the other hand, relies on entrusting other people to share the burden of understanding, a topic that seems underexplored in current XAI research considering its importance in understanding. We argue that future work should examine group settings and verbal explanation modalities for possibilities of including outsourcing as an understanding strategy.
5.2. Using explanatory stances to address different aims of affected stakeholders
In this section, we discuss how future explanation approaches can incorporate the stakeholders’ aims and explanation’s purpose in the design process by using the concept of explanatory stances (Keil, 2021, 2006; Lombrozo and Wilkenfeld, 2019). We first discuss how participants’ aims emerged in our study and second outline how explanatory stances can be used to convey information according to these aims.
5.2.1. Considering the stakeholder’s aim and the explanation’s purpose.
Explanations have the purpose of enabling stakeholders to attain their aims (Freiesleben and König, 2023), such as calibrating their acceptance or trust towards a system (Shin, 2021), or performing an action, like contesting a decision or requesting human intervention (Alfrink et al., 2022). Participants in our study were asked to vote over the adoption of the ADM systems to create a purpose which would incentivize acquisition of information through question-driven explanations. This worked for participants who used the explanations to underpin their decision and in turn reported a slight increase in decision confidence (Table 3).
On the other hand, the perceived risks and benefits of the ADM systems seems to have influenced how participants defined their aims and approached inquiry: Participants inquired more about societal and structural factors in the employment use case, and more about practical and operational factors in the wristband use case. Explanation approaches for affected stakeholders should thus consider that the domain of the use case and stakeholders’ perceptions towards it impact which information is deemed relevant. Acknowledging the large role that the institutional context plays in how affected stakeholders perceive ADM (Brown et al., 2019; Woodruff et al., 2018; Lee et al., 2019; de Fine Licht and de Fine Licht, 2020; Veale et al., 2018), explanations likely cannot always meet the aims of all stakeholder groups, especially when conflicts of interest exist.
5.2.2. Explanatory stances can be useful to guide information acquisition.
@@ 5.2. Using explanatory stances to address different aims of affected stakeholders
People can differ considerably in how they approach information acquisition and understanding, depending in part on their stakeholder role and their aims. As explanatory stances (mechanical, design, and intentional (Dennett, 1998)) describe the ways in which people make sense of information and generalize from it, applying them in explanation design could help to convey information more effectively. Previous XAI research considered explanatory stances in conceptual approaches (Páez, 2019; Miller, 2019; Zerilli, 2022), but rarely empirically. We argue that future work in XAI should investigate explanatory stances for two main reasons:
First, our interviews showed that certain topics lend themselves for the application of certain stances. For example, participant inquiries about system specifications mostly applied a mechanical or design stance, while inquiries about system context rather applied the design and intentional stance (Section 4.2.2). Which stance is applied can depend on the stakeholder role, for example: Features of an ADM system can be relevant for both developers, who could inquire about the features’ data types, and affected stakeholders, who could inquire how features capture a real world context. Importantly, both stakeholders could ask the same question: What features does the system consider and why? Considering which stance a stakeholder likely applies in acquiring information might therefore allow for better personalization of question-driven explanations.
@@
Second, encouraging application of different explanatory stances can support a change in perspective. Work in the cognitive sciences describes the possibility that the design stance might be preferred over the mechanical stance by default (Lombrozo and Wilkenfeld, 2019).111111Preference in explanatory stances is a controversial discussion, the literature currently does not draw general conclusions (Keil, 2021; Lombrozo and Wilkenfeld, 2019; Kelemen et al., 2013). In our study, we observe that both the design and intentional stance in fact emerged more frequently than the mechanical stance in the first inquiry phase. Therefore, explanation approaches can encourage an active change in perspective by providing material that uses a different stance than the participant and steers attention to information that is not explored autonomously. For example, introducing the XAI Question Bank (Liao et al., 2020) in the second inquiry phase of our study incited participants to apply a mechanical stance, which they did less in the first (unguided) phase of inquiry.
5.3. Taking steps to incorporate affected stakeholders into human-centered XAI
Affected stakeholders, as decision subjects, are arguably the most vulnerable stakeholder group of ADM and should therefore be able to receive explanations about systems that might affect them (European Parliament. Directorate General for Parliamentary Research Services., 2019; Commission, 2021; Vredenburgh, 2022; Panigutti et al., 2023). Our findings draw attention to two challenges in using question-driven explanations for this purpose: First, participants in our study expressed interest to learn about intentional information, which is not typically covered in explanations. Second, affected stakeholders might have no interest in interacting with explanations or in taking part in decisions about ADM. In the following, we briefly discuss these issues with a perspective on future work in XAI research.
5.3.1. Handling explanations of intention.
In our study, many participants applied the ”intentional stance” (Dennett, 1998) to inquire about the ADM systems’ context and usage, aiming to examine the beliefs and values ingrained in the systems (Section 4.2). These inquiries were often motivated by how participants perceived the deploying institution and the involved risks and benefits (Brown et al., 2019; Woodruff et al., 2018). However, explaining the intentions and values of a system’s deployment means taking a stance towards the values involved. Further, intentional information is not strictly factual and thus might not be easily accessible. Then again, human explanations are by nature social (Miller, 2019) and transactional (Keil, 2006), and tend to use subjective views to better convey information.
Consequently, determining the explanation’s purpose and stakeholders’ aim might facilitate the decision of whether intentional information should be included (Freiesleben and König, 2023). For example, if the explanation’s purpose is a solely mechanical or functional description of the system in order to enable operation, intentional information is likely negligible. In contrast, when stakeholders should form value assessments or investigate possibilities for contestation (Alfrink et al., 2022), intentional explanations can contribute meaningful information, e.g., by using ”justifications” (Biran and Cotton, 2017; Biran and McKeown, 2017; de Fine Licht and de Fine Licht, 2020) to explain on which grounds decisions are made. As stakeholders’ interest in the beliefs and values of ADM is a central information need, we argue that future research should examine how this information can be conveyed meaningfully.
5.3.2. Taking steps to connect explainability to civic education.
Our studies show that affected stakeholders are not always interested in understanding ADM, even if it might affect them. Assuming that one goal of explainable AI is to provide the most information to the most people in order to increase understanding of public AI systems, the question of how to redeem the interest of these stakeholders is central. Work on similar issues in democratic participation terms this issue ”civic ignorance”: people’s general disinterest in laws and regulations that shape their lives (Lau and Redlawsk, 2001; Lupia, 2016). In fact, civic ignorance is not meant as a stigma, but as a common effect that applies to most people, including political experts (Lupia, 2016). The corresponding research describes viable remedies, beginning with identifying the intersection of i) information that the public is interested in because it advances their core concerns and ii) information that is sufficient to enable performance of a task, such as judging a political topic. When this intersection, the ”information that matters” (Lupia, 2016), has been outlined, subsequent steps can develop educational measures that have better chances to redeem public interest.
Regarding XAI, this paper aims to accomplish the first step of identifying information that is both relevant to affected stakeholders and sufficient to form a judgment. Recent work in AI literacy has developed catalogs of competences and design considerations for educational interventions on AI (Long and Magerko, 2020; Ng et al., 2021). We argue that future work should combine these empirical and conceptual insights to design explanations that effectively convey information to affected stakeholders.
5.4. Implications
@@
In this section, we briefly summarize the implications that our findings and discussion have for the design of explanations for affected stakeholders.
Affected stakeholders’ information needs are diverse and differ between use cases. Our proposed XAI Novice Question Bank (Figure 2) extends the well-known XAI Question Bank (Liao et al., 2020) by including questions on the system context and usage. It can thus be leveraged to design better-targeted explanations for affected stakeholders.
Affected stakeholders’ information needs depend on the actions they want to take, such as assessing values, voting, or contestation. Explanations should support affected stakeholders in realizing their specific aim(s).
Question-driven explanations allow stakeholders to decide when their understanding is sufficient for their aims. But assessing one’s own understanding is inherently difficult and prone to cognitive biases (Section 5.2). Facilitating this assessment could be necessary to avoid the illusion of explanatory depth in question-driven explanations.
Information can be understood from different angles. Certain angles might be more intuitive for certain stakeholders: domain experts might ask about a feature’s purpose (functional) while job-seekers might ask why it was included (intentional). The chosen angle can impact which explanation is suitable.
@@
Information on the underlying reasoning of why an ADM system is deployed can be relevant for affected stakeholders. Typical explanations focus on the workings of the system rather than this type of contextual information. XAI research should also provide guidelines on how the decision-making behind a system’s development and deployment can be explained.
@@
Researchers developing explanations for affected stakeholders ought to be aware that people might be disinterested in explanations despite being affected, caused, for example, by a perceived inability to understand ADM or a feeling of powerlessness. Future research should examine further how these factors impact explanations for affected stakeholders.

### v6

@@ 5. Discussion
This section outlines the theoretical and practical implications of our study. It addresses AI novices’ information needs, the interaction between question-driven explanations and understanding, and the inclusion of domain experts and decision subjects in explanation design. These implications are summarized at the end (Section 5.3).
5.1. Theoretical implications
5.1.1. Designing explanations for AI novices
@@
One of XAI’s enduring key questions is: What makes a “good” explanation? Recent work outlines that, to produce understanding, explanations should be contrastive, selected, and social while being first and foremost contextual (Miller, 2019), that they should interface directly with people’s causal reasoning (Byrne, 2023) and consider the learner’s aim and prior knowledge (Mueller et al., 2019). Fulfilling these requirements puts a high demand on explanation design. Further, existing explanation approaches do not typically address the needs of AI novices (Shin, 2023), even though they compose multiple stakeholder groups, including people affected by algorithmic decisions and end-users. We thus argue that one step in designing explanations for AI novices is to identify their information needs and examine how they depend on participants’ perceptions, stakeholder roles, and domain expertise. In the following, we present how the findings of our study support this design approach.
@@
Our findings show that AI novices’ information needs cover all parts of an ADM system’s “lifecycle” (Dhanorkar et al., 2021), from technical details over deployment consequences to different dimensions of usage. Our proposed “XAI Novice Question Bank” (Figure 3) provides a summary of these topics, which can inform future explanation design. However, we also observe differences in information needs between use cases and differences to Liao et al. (2020)’s study with AI practitioners. This is in line with previous work, which shows that various contextual factors mediate which information needs participants articulate, including familiarity with the presented system (Kramer et al., 2018), tangibility of the system (Long et al., 2021), perceived risks and benefits (Araujo et al., 2020) (e.g., trade-off between safety and privacy in the wristband use case), domain expertise (Wang and Yin, 2021), and technical expertise (Cheng et al., 2019).
5.1.2. Anticipating and handling explanations of intention.
Explanations have the purpose of enabling stakeholders to attain their aims (Langer et al., 2021), such as facilitating the evaluation of an ADM system’s values (Shin, 2021, 2023), or performing an action, like contesting a decision or requesting human intervention (Alfrink et al., 2023). In our study, participants were asked to vote over adopting the ADM systems to incentivize information acquisition through questions. I.e., the “purpose” (Freiesleben and König, 2023) of the explanations was to enable a well-informed decision. For several participants, this succeeded, as they used the explanations to underpin their decision and reported a slight increase in decision confidence (Table 3).
At the same time, as participants’ information needs differed between use cases, their perception of what a well-informed decision would need to consider differed. Participants tended to inquire more about societal and structural factors in the employment prediction use case and more about practical and operational aspects in the wristband use case. Framed in the context of the explanatory stances, many participants used the “intentional stance” (Dennett, 1998) to ask questions about the AMS algorithm and the beliefs and values it represents (Section 4.2). This demand for explanations of intention was likely due to the system’s close association with the job agency, as previous work has shown that the reputation of the deploying institution can affect perceptions of ADM systems (Brown et al., 2019; Woodruff et al., 2018; Lee et al., 2019).
However, delivering explanations of the intention behind a system’s deployment is difficult, as it means interpreting the values involved and choosing how to present them – a process that rather should be performed by each participant individually due to its inherent subjectivity. Further, information about the true purpose of an ADM system’s deployment can be hard to attain, as it is often part of a deploying institution’s non-public decision-making process and can be kept intentionally opaque (Rudin, 2019).
Consequently, whether explanations of intention should be included needs to be matched to the explanation’s purpose. For example, if the explanation’s purpose is solely a mechanical or functional description of the system to enable operation, intentional information is likely negligible. In contrast, when stakeholders should form value assessments or investigate possibilities for contestation (Alfrink et al., 2023), intentional explanations can contribute meaningful information, e.g., by using “justifications” (Biran and Cotton, 2017; Biran and McKeown, 2017; de Fine Licht and de Fine Licht, 2020) to explain how decisions are made. As stakeholders’ interest in intentional information was central to our study, future research should examine how this information can be conveyed meaningfully.
5.1.3. Taking steps to connect explainability to civic education.
Our studies show that our participants were not always interested in understanding ADM systems, even though they might affect them. Assuming that democratic civil societies have the right to understand public AI systems (Züger and Asghari, 2023) and that explanations are a means to realize this right (de Fine Licht and de Fine Licht, 2020), the question of how to redeem the interest of these stakeholders becomes essential. Research on democratic participation uses the term “civic ignorance”, which is the common phenomenon that citizens are disinterested in laws and regulations shaping their lives (Lau and Redlawsk, 2001; Lupia, 2016). Civic ignorance is not a stigma but an effect that applies to most people, including political experts (Lupia, 2016). Lupia (2016) describes that interest can be remedied by developing educational material that covers the “information that matters”, which is the intersection of i) information that advances the core concerns of the public and ii) information that enables the performance of a task, such as judging a political topic.
This paper aims to accomplish this step of identifying information relevant to people affected by algorithmic decisions and sufficient to form a judgment. We argue that future work should combine these empirical insights with conceptual work on AI literacy (Long and Magerko, 2020; Ng et al., 2021) to design explanations that effectively convey information to these stakeholder groups. Our study outlines the first steps towards this approach: Orientating explanations along AI novices’ information needs, allowing for flexible switching between different topics and explanatory stances, using guidance material such as the XAI Question Bank (Liao et al., 2020) and the XAI Novice Question Bank (Figure 3), and considering to include contextual information such as intention.
5.2. Practical implications
@@
The direct aim of explanations is to increase a person’s understanding of an ADM system (Langer et al., 2021), but both designing explanations that reliably increase understanding and evaluating if they succeeded is surprisingly difficult (Schmude et al., 2023). In the following, we analyze how our question-driven explanations interacted with participant understanding and derive two concrete implications for future explanation design.
First, question-driven explanations succeed in letting participants independently select relevant information and guiding them to non-explored aspects of the ADM system, for example, by providing the XAI Question Bank (Liao et al., 2020) as guidance. This supports the notion of personalized explanations (Conati et al., 2021; Shulner-Tal et al., 2022; Martijn et al., 2022) and could be further elaborated in a design study of a digital question-driven explanation tool based on the XAI Novice Question Bank’s categories. This would allow for a better overview and more straightforward navigation of information while potentially being suitable and relevant for non-expert users. This implementation could further explore the possibility of including a generative model using an information base to generate explanations in response to users’ input queries, as has already been proposed for other domains (Zhao et al., 2024).
While this generative explainer would allow for flexible and comprehensive explanations that could vary in detail, it also poses several challenges: First, the information used for our explanations was “scavenged” (Wieringa, 2023) from various sources, meaning that it needs to be decided which information should be prioritized if sources are ambiguous or contradict each other. Second, non-factual information, such as explanations of intention, would require the model to reflect multiple points of view clearly and neutrally represent value conflicts likely to arise in high-risk ADM systems (Raji et al., 2022). Third, non-deterministic generative models are known to produce content that can vary in quality and might even contain wrong information (Karinshak et al., 2023), which would be very difficult to spot for AI novices. Lastly, multiple participants in our study reported that they appreciated the social nature of the explanation format and the exchange with another person, which is in line with previous research (Miller, 2019) and which would be mostly lost when transferred to a generative explainer. Providing a viable implementation of generative question-driven explanations would thus require thoroughly considering these issues.
@@
Second, several processes can interfere with increasing understanding through question-driven explanation, including blind spots, outsourcing, and understanding self-assessment. We use the term blind spots for information that participants felt was missing but that they could not acquire. This might mean a lack of “intelligibility types” (Lim and Dey, 2009), i.e., a lack of guidance on which information was available and how to ask about it. Documents such as our proposed XAI Novice Question Bank (Figure 3) and the XAI Question Bank (Liao et al., 2020) should help in addressing this issue by showing people the relevant categories of information.
Outsourcing, on the other hand, refers to sharing the process of understanding with other people to make sense of it collaboratively, a topic that seems underexplored in current XAI research. Future work should thus consider if the effectiveness of explanations can be increased by providing them to groups of people who can rely on the “expertise in other minds” (Keil, 2006), i.e., discuss and collaborate. An essential aspect of outsourcing understanding is the reliance on other people to aid in the understanding process (Keil, 2006), especially when the topic seems complex and overloaded. In our study, this prompted participants to skip parts of the understanding process entirely and instead ask the study examiner for their assessment. Considering that civil society organizations are already entrusted with providing support for people regarding finances, education, legal aid, and more, they are an intuitive point of orientation when it comes to providing explanations about high-risk ADM systems to those affected by their decisions (Scott et al., 2022). We argue that future work in XAI should thus cooperate with these organizations to deploy explanations that enable AI novices to understand, negotiate, and contest decisions of ADM systems (Kaun and Taranu, 2020).
Lastly, question-driven explanations also require understanding self-
assessment, meaning people decide when their understanding is sufficiently developed. However, this can be difficult to assess (Keil, 2006), as perceived understanding can paradoxically decrease despite a gain in information (Cheng et al., 2019) and can also fall prone to the “illusion of explanatory depth” (Rozenblit and Keil, 2002; Chromik et al., 2021), an overestimation of one’s understanding. Previous work has outlined various strategies to offset these issues, including deliberate self-explanation (Duckworth, 2001), better initial introductions to the system (Chromik et al., 2021), and checking for people’s factual understanding by asking examination questions (Buçinca et al., 2020). To offset the decrease in self-reported understanding stemming from the gain in new information, we propose to include an elicitation of participants’ information gain in addition to their perceived understanding. A survey question could be formulated as “How much knowledge did you gain about the system?”.
5.2.2. Matching explanations to stakeholders’ needs.
@@
For the reasons outlined, people can take very different approaches to information acquisition and understanding, depending partly on their stakeholder role and aims. As explanatory stances (mechanical, design, and intentional (Dennett, 1998)) describe how people make sense of information and build predictive strategies from it, they can inform explanation design by considering which stance people apply when acquiring information. While previous XAI research explored explanatory stances conceptually (Páez, 2019; Miller, 2019; Zerilli, 2022), we argue that future work should investigate explanatory stances empirically for two main reasons:
First, our analysis shows that participants tended to apply different stances in their questions depending on the topic of inquiry: questions about data and system details tended to aim for technical information (mechanical), while questions about usage and context instead aimed for information on the system’s practical and conceptual design aspects (design) or the values that it might enforce due to how and why it was developed (intentional). Importantly, participants can assume any of the three stances in their questions, and similar-sounding questions might aim for different information, such as “How often does the system make mistakes?”. Identifying a question’s explanatory stance can thus help match the provided information to the person’s information need by considering the question’s context or disambiguating which category of information it aims for. This identification can be facilitated by context information, such as the participants’ previous questions, or by explicitly inquiring how the information is relevant to the participant.
@@
Second, encouraging the use of different explanatory stances can support a change in perspective. Work in the cognitive sciences, for example, describes the possibility that the design stance might be preferred over the mechanical stance by default (Lombrozo and Wilkenfeld, 2019).121212Preference in explanatory stances is a controversial discussion (Keil, 2021; Lombrozo and Wilkenfeld, 2019; Kelemen et al., 2013). In our study, we observed that participants used the design and intentional stance instead of the mechanical stance, especially before receiving the XAI Question Bank (Liao et al., 2020). However, the boundaries might not be as clear-cut due to the interpretation required when assigning explanatory stances to participants’ questions. Still, considering which information might prompt participants to take a specific stance could be useful for the design of explanations, for example, by providing material that prompts a variety of stances and steers attention to information that is not explored autonomously. For example, introducing the XAI Question Bank (Liao et al., 2020) in the second inquiry phase of our study encouraged participants to ask questions aiming for technical details (mechanical), which they reported to be very useful as they had not considered these perspectives before.
5.3. Summary of implications
@@
In this section, we briefly summarize the implications that our findings and discussion have for the design of explanations.
Theoretical
AI novices’ information needs vary by use case. The proposed XAI Novice Question Bank (Figure 3) expands the XAI Question Bank (Liao et al., 2020) with questions on system context and usage, highlighting relevant information for novices. Future work should explore integrating this information into formats that can be flexibly adapted to different systems.
@@
Information on why an ADM system is deployed can be relevant for various stakeholders. Typical explanations focus on the system’s workings but not on this contextual information. XAI research should consider how to best convey information on the intention behind a system’s development and deployment.
@@
Researchers developing explanations for those affected by algorithmic decisions should note that people might be disinterested in explanations due to a perceived inability to understand ADM or feelings of powerlessness. Future research should further examine how these factors impact explanations for groups who are affected by ADM systems.
Practical
Question-driven explanations let stakeholders acquire information until their understanding meets their aims. A digital implementation with a generative explainer model could apply these explanations practically. Challenges to understanding can be addressed by guiding available information, considering if understanding can be outsourced to others, and eliciting information gain in addition to perceived understanding.
Information can be understood from different angles, and some may be more intuitive for specific stakeholders: domain experts might ask about a system’s design aspects. In contrast, job-seekers might ask about the beliefs governing its deployment. Explanatory stances can help to identify which angle someone will take, allowing for better matching of explanations to their needs. Encouraging stance-switching, such as providing technical information to AI novices, can lead to valuable perspective changes.